Two-tailed test, o? is known, a= 0.05, n= 45


1
Expert's answer
2022-05-03T16:20:10-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=\mu_0"

"H_1:\mu\not=\mu_0"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}

."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu_0}{\sigma/\sqrt{n}}"